# Deep polytopic autoencoders for low-dimensional linear parameter-varying approximations and nonlinear feedback design

### v1

@@ 4 Higher order series expansions of LPV approximations in SDRE feedback design
For the quasi-LPV system eq. 2, the nonlinear SDRE feedback design defines the input $u$ as
$u(t)=-B^{\mkern-1.5mu\mathsf{T}}P(\rho(v(t)))v(t),$ (8)
@@
Since the repeated solve of eq. 9 is costly, series expansions of the feedback law eq. 8 have been proposed.
We briefly derive the formulas for higher order expansions with nonlinear parameter dependencies in $A$. For linear parameter dependencies, the first-order approximation has been developed in [5] for a single parameter dependency and has been extended in [1] to the multivariate case.
@@
$P(\rho)\approx P_{0}+\sum_{1\leq\lvert\alpha\rvert\leq p}\rho^{(\alpha)}P_{\alpha},$
where $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbb{N}^{r}$ is a multiindex with $\lvert\alpha\rvert:=\sum_{i=1}^{r}\alpha_{i}$, where $\rho^{(\alpha)}:=\rho_{1}^{\alpha_{1}}\rho_{2}^{\alpha_{2}}\cdots\rho_{r}^{\alpha_{r}}$, and where, importantly, $P_{\alpha}$ are constant; see [16] for more details.
For nonlinear parameter dependencies, we will consider a truncated series expansion of $A$:
@@
where $A_{\alpha}$ are constant matrices; see eq. 6 that describes the first and second order expansion in the considered case with the structure of the nonlinear decoding as $\rho\to W(c(\rho)\otimes\rho)$.
We insert the expansions of $A$ and $P$ into the SDRE eq. 9 and solve for the coefficients $P_{\alpha}$ by matching the individual powers of $\rho$. For $\alpha=0$, we obtain the standard Riccati equation
@@
whereas for $\lvert\alpha\rvert=1$, the coefficients $P_{\alpha}$ are obtained from
$\begin{split}P_{0}\bigl{(}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\bigr{)}+\bigl{(}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\bigr{)}^{\mkern-1.5mu\mathsf{T}}P_{0}\\ {}-{}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}BB^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}BB^{\mkern-1.5mu\mathsf{T}}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}=0\end{split}$
for arbitrary values of $\rho$. This amounts for the $P_{\alpha}$ to be given via the Lyapunov equations
$\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)^{\mkern-1.5mu\mathsf{T}}P_{\alpha}+P_{\alpha}\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)$
$=-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha},$ (12)
for the $r$ multiindices in $\alpha\in\mathbb{N}^{r}$ for which $\lvert\alpha\rvert=1$; see also [1, 16] where the same formulas have been derived. Similarly, we may collect the terms in which $\rho$ appears quadratic. Let ${\alpha^{*}}$ be a multiindex with $\lvert{\alpha^{*}}\rvert=2$ and let ${\beta^{*}}$ and ${\delta^{*}}$ be those two multiindices with $\lvert{\beta^{*}}\rvert=\lvert{\delta^{*}}\rvert=1$ that satisfy ${\beta^{*}}+{\delta^{*}}={\alpha^{*}}$. Then, the coefficient $P_{{\alpha^{*}}}$ that cancels the corresponding contributions in the equation is defined through the Lyapunov equation
$(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})^{\mkern-1.5mu\mathsf{T}}P_{{\alpha^{*}}}+P_{{\alpha^{*}}}(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})$
@@
$\quad{}-{}A_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\beta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\delta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}.$ (13)
Note that for given $r$ and $p$, there exist exactly $\binom{r+p-1}{p}$ unique such multiindices $\alpha$ with the absolut value of $\rvert\alpha\lvert=p$. For such a series expansion of $P$, the feedback law eq. 8 reads
$u(t)=-B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{\rvert\alpha\lvert\leq p}\rho^{(\alpha)}P_{\alpha}\bigr{)}v(t)=:u_{P_{|\alpha|\leq p}}(t),$ (14)
which with the precomputed $B^{\mkern-1.5mu\mathsf{T}}P_{\alpha}$ is efficiently evaluated by adapting the values of $\rho$. We note that for $p=0$ this formula reduces to the standard linear quadratic regulator (LQR) feedback with respect to a linearization about $v^{*}=0$.
@@ 5 Solving the high-dimensional matrix equations
To solve the occurring high-dimensional linear and nonlinear matrix equations eqs. 11, 12 and 13, some reformulations and state-of-the-art approaches are needed. Assume that the dimensions of the input and output operators $m,\ell$ are much smaller than the state-space dimension $n$ and that $A_{0}$ in eq. 10 is sparsely populated. The first equation eq. 11 is the standard factorized Riccati equation, which under the assumptions above can be efficiently solved for a low-dimensional approximation of the solution $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, with $Z_{0}\in\mathbb{R}^{n\times k}$ and $k\ll n$, e.g., via the low-rank Newton-Kleinman iteration [7]. For the two types of Lyapunov equations eqs. 12 and 13, some reformulations are needed to efficiently compute the solutions in the large-scale sparse setting; namely, the right-hand sides need to be reformulated into symmetric low-rank factorizations. In the case of eq. 12, given the approximation $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, the right-hand side can be rewritten as
$-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha}\approx-\begin{bmatrix}A_{\alpha}^{\mkern-1.5mu\mathsf{T}}Z_{0}&Z_{0}\end{bmatrix}J\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{\alpha}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$
@@
where $J=\begin{bmatrix}0&I_{k}\\ I_{k}&0\end{bmatrix}$. For Lyapunov equations eq. 12 with such symmetric factorized right-hand sides, the ADI approach has been extended to compute appropriate solution approximations of the form $L_{\alpha}D_{\alpha}L_{\alpha}^{\mkern-1.5mu\mathsf{T}}\approx P_{\alpha}$; see [18]. Similarly, we can derive for the right-hand side of eq. 13 low-rank factorizations of the form $LDL^{\mkern-1.5mu\mathsf{T}}$, where
$L^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{{\alpha^{*}}}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\delta^{*}}}-BK_{{\delta^{*}}})\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\beta^{*}}}-BK_{{\beta^{*}}})\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},D=\begin{bmatrix}J&0&0\\ 0&0&D_{{\beta^{*}},{\delta^{*}}}\\ 0&D_{{\beta^{*}},{\delta^{*}}}&0\end{bmatrix},$
with $D_{{\beta^{*}},{\delta^{*}}}=\begin{bmatrix}D_{{\beta^{*}}}&0\\ 0&D_{{\delta^{*}}}\end{bmatrix}$, the low-rank factorizations $L_{{\beta^{*}}}D_{{\beta^{*}}}L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\beta^{*}}}$ and $L_{{\delta^{*}}}D_{{\delta^{*}}}L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\delta^{*}}}$, and the feedback matrices $K_{{\beta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}$ and $K_{{\delta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}$.

### v2

@@ 4 Higher order series expansions of LPV approximations in SDRE feedback design
Although, as shown in Lemma 1, the second-order expansion of the coefficient function $A$ coincides with the expansion of first order, the feedback law will still exhibit second-order dependencies as we will lay out in this section. For the quasi-LPV system eq. 2, the nonlinear SDRE feedback design defines the input $u$ as
$u(t)=-B^{\mkern-1.5mu\mathsf{T}}P(\rho(x(t)))x(t),$ (10)
@@
We note that for quasi-LPV systems the parameter changes continuously with the states. Since the repeated solve of eq. 11 is costly, series expansions of the feedback law eq. 10 have been proposed. We now briefly derive the formulas for higher order expansions of the solution to eq. 11 with nonlinear parameter dependencies in $A$. For linear parameter dependencies, the first-order approximation has been developed in [10] for a single parameter dependency and has been extended in [1] to the multivariate case.
@@
$P(\rho)\approx P_{0}+\sum_{1\leq\lvert\alpha\rvert\leq p}\rho^{(\alpha)}P_{\alpha},$ (12)
where $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbb{N}^{r}$ is a multiindex with $\lvert\alpha\rvert:=\sum_{i=1}^{r}\alpha_{i}$, where the coefficients are $\rho^{(\alpha)}:=\rho_{1}^{\alpha_{1}}\rho_{2}^{\alpha_{2}}\cdots\rho_{r}^{\alpha_{r}}$, and where, importantly, all $P_{\alpha}$ are constant symmetric matrices, namely
$P_{\alpha}=\frac{1}{\alpha_{1}!\alpha_{2}!\cdots\alpha_{r}!}\frac{\partial^{\lvert\alpha\rvert}}{\partial_{\rho_{1}}^{\alpha_{1}}\partial_{\rho_{2}}^{\alpha_{2}}\cdots\partial_{\rho_{r}}^{\alpha_{r}}}P(0).$
We note that, because the function $P$ is not known, this formula is merely stated to confirm that the $P_{\alpha}$ are constant matrices. Practically, in what follows, we will compute the $P_{\alpha}$ directly by equating coefficients in truncated series.
For the nonlinear parameter dependencies in $A(\rho)$, we will consider a truncated series expansion of $A$:
@@
where $A_{\alpha}$ are constant matrices; see eq. 9 that describes the first and second-order expansions as follows: For $p=1$, we can identify all $\lvert\alpha\rvert=p$ with $k=1,\dotsc,r$ and obtain $\rho^{(\alpha)}A_{\alpha}=\rho_{k}\widetilde{A}(w_{k})$, where $w_{k}$ is the $k$-th column of $W$. Following Lemma 1, in the considered case with the structure of the nonlinear decoding as $\rho\to W(c(\rho)\otimes\rho)$, the second-order expansion coincides with the first-order one. We also note that, if $\widetilde{A}$ is known (see Remark 1), then $A_{0}$ is readily inferred without knowledge of $\rho$ or $W$.
We insert the expansions of $A$ and $P$, namely eqs. 12 and 13, into the SDRE eq. 11 and solve for the coefficients $P_{\alpha}$ by matching the individual powers of $\rho$. For $\alpha=0$, we obtain the classical standard Riccati equation
@@
as it is used for the design of feedback controllers for linear systems. Moving on, for $\lvert\alpha\rvert=1$, the coefficients $P_{\alpha}$ are obtained from requiring that the first-order terms fulfill
$\begin{split}P_{0}\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\right)+\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\right)^{\mkern-1.5mu\mathsf{T}}P_{0}-\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}BB^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}BB^{\mkern-1.5mu\mathsf{T}}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}=0,\end{split}$
for arbitrary values of $\rho$. This defines the $P_{\alpha}$ as the solutions to the Lyapunov equations
$\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)^{\mkern-1.5mu\mathsf{T}}P_{\alpha}+P_{\alpha}\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)=-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha},$ (15)
for the $r$ multiindices in $\alpha\in\mathbb{N}^{r}$ for which $\lvert\alpha\rvert=1$ holds; see also [1, 30] where the same formulas have been derived. Similarly, we may obtain the $P_{\alpha}$ by setting the terms in which $\rho$ appears quadratic to zero:
00 0 $=P_{0}\left(\sum_{\lvert\alpha\rvert=2}\rho^{(\alpha)}A_{\alpha}\right)+\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}\right)\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\right)+\left(\sum_{\lvert\alpha\rvert=2}\rho^{(\alpha)}P_{\alpha}\right)A_{0}+\dotsc$
$\qquad{}-{}\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}\right)BB^{\mkern-1.5mu\mathsf{T}}\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}\right)-\left(\sum_{\lvert\alpha\rvert=2}\rho^{(\alpha)}P_{\alpha}\right)BB^{\mkern-1.5mu\mathsf{T}}P_{0}-\dotsc,$
[…]
$(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})^{\mkern-1.5mu\mathsf{T}}P_{{\alpha^{*}}}+P_{{\alpha^{*}}}(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})$ (16)
@@
$\quad{}-{}A_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\beta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\delta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}.$
Note that for given $r$ and $p$, there exist exactly $\binom{r+p-1}{p}$ unique such multiindices $\alpha$ such that $\rvert\alpha\lvert=p$.
Finally, we note that for such a series expansion of $P$, the feedback law eq. 10 reads
$u(t)=-B^{\mkern-1.5mu\mathsf{T}}\left(\sum_{\rvert\alpha\lvert\leq p}\rho^{(\alpha)}P_{\alpha}\right)x(t)=:u_{P_{\lvert\alpha\rvert\leq p}}(t),$ (17)
which with the precomputed $B^{\mkern-1.5mu\mathsf{T}}P_{\alpha}$ is efficiently evaluated by adapting the values of $\rho$. We also note that for $p=0$, this formula reduces to the standard linear quadratic regulator (LQR) feedback with respect to a linearization about $x^{\ast}=0$.
@@ 5 Solving the high-dimensional matrix equations
To compute the solutions to the occurring high-dimensional matrix equations eqs. 14, 15 and 16, some reformulations and state-of-the-art approaches are needed. In the considered setup, we can exploit that the ranks of $B$ and $C$ are much smaller than the state-space dimension $n$ and that $A_{0}$ in eq. 13 is sparsely populated, so that the first equation eq. 14—a standard Riccati equation—can be efficiently solved for a low-rank approximation of the symmetric positive semi-definite stabilizing solution $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, with $Z_{0}\in\mathbb{R}^{n\times k}$ and $k\ll n$; see [11, 42], e.g., via the low-rank Newton-Kleinman iteration [14, 39], the incremental subspace iteration [36] or the RADI method [12, 15].
However, for the two types of Lyapunov equations eq. 15 and eq. 16 occurring through the series expansion, some reformulations are needed to compute the solutions in this large-scale setting; namely, the right-hand sides need to be reformulated into symmetric low-rank factorizations. Given the approximation $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$ of the initial Riccati equation eq. 14, the right-hand side in eq. 15 can be rewritten as
$-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha}\approx-\begin{bmatrix}A_{\alpha}^{\mkern-1.5mu\mathsf{T}}Z_{0}&Z_{0}\end{bmatrix}J\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{\alpha}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$ (18)
@@
where the symmetric yet indefinite center matrix is given by $J=\begin{bmatrix}0&I_{k}\\ I_{k}&0\end{bmatrix}$, with $I_{k}$ denoting the $k$-dimensional identity matrix and $k$ being the number of columns of $Z_{0}$. For Lyapunov equations eq. 15 with such symmetric factorized right-hand sides eq. 18, the ADI approach has been extended to compute appropriate solution approximations of the form $L_{\alpha}D_{\alpha}L_{\alpha}^{\mkern-1.5mu\mathsf{T}}\approx P_{\alpha}$ following the format of the right-hand side; see [35]. Furthermore, while the coefficient matrix in eq. 15, namely
$A-BB^{\mkern-1.5mu\mathsf{T}}P_{0}=A-BV^{\mkern-1.5mu\mathsf{T}}$ (19)
with $V=B^{\mkern-1.5mu\mathsf{T}}P_{0}\in\mathbb{R}^{n\times m}$, is dense and high-dimensional, it is not necessary to assemble this matrix at any point because eq. 19 is a large-scale sparse matrix with low-rank perturbation. Solution methods for large-scale Lyapunov equations only rely on matrix-vector operations or the solution to linear systems with eq. 19. The latter can be efficiently computed using the Sherman-Morrison-Woodbury matrix inversion formula or augmented matrix approach [27] such that only the solution of linear systems with large-scale sparse coefficient matrices and matrix-vector products are needed. It is important to observe that the coefficient eq. 19 is Hurwitz, i.e., all eigenvalues lie in the left open half-plane, such that the established solution methods can be applied.
Similarly, we can derive for the right-hand side of eq. 16 symmetric indefinite low-rank factorizations of the form $LDL^{\mkern-1.5mu\mathsf{T}}$. Thereby, the factor are
$L^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{{\alpha^{*}}}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\delta^{*}}}-BK_{{\delta^{*}}})\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\beta^{*}}}-BK_{{\beta^{*}}})\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$ (20)
and
$D=\begin{bmatrix}J&0&0&0&0\\ 0&0&0&D_{{\beta^{*}}}&0\\ 0&0&0&0&D_{{\delta^{*}}}\\ 0&D_{{\beta^{*}}}&0&0&0\\ 0&0&D_{{\delta^{*}}}&0&0\end{bmatrix},$ (21)
where we have the approximations of the solutions to the Lyapunov equations eq. 15 as $L_{{\beta^{*}}}D_{{\beta^{*}}}L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\beta^{*}}}$ and $L_{{\delta^{*}}}D_{{\delta^{*}}}L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\delta^{*}}}$, and the intermediate feedback matrices $K_{{\beta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}$ and $K_{{\delta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}$. As for eq. 15, this reformulation of the constant terms of eq. 16 allows the application of the same computational procedures for the solution, for example, the $LDL^{\mkern-1.5mu\mathsf{T}}$-factorized ADI method [35]. The coefficient matrix in eq. 16 is the same as in eq. 15 such that it has the same properties being sparse plus low-rank and Hurwitz. Therefore, we can apply the same numerical approaches to solve eq. 16 as for eq. 15.
In both cases of Lyapunov equations eqs. 15 and 16, we have that the inner dimensions of the reformulated right-hand side terms eqs. 18, 20 and 21 depend on dimensions of the factors of the solution approximations to eqs. 14 and 15. Therefore, we can expect the dimensions of the right-hand side factors to be larger than the dimensions of the input and output matrices used in the Riccati equation eq. 14. In general, we strongly recommend to compress the right-hand side factors at least to their numerical rank using economy-size QR and eigenvalue decompositions. All right-hand sides of Lyapunov equations considered in this paper can be written as indefinite symmetric factorizations of the form $LDL^{\mkern-1.5mu\mathsf{T}}$, with $L\in\mathbb{R}^{n\times k}$ and $D\in\mathbb{R}^{k\times k}$ symmetric. For the compression, first, compute the economy-size QR of $L=QR$, with $Q\in\mathbb{R}^{n\times k}$ and $R\in\mathbb{R}^{k\times k}$. Next, compute the eigenvalue decomposition of the symmetric matrix
$RDR^{\mkern-1.5mu\mathsf{T}}=U\Lambda U^{\mkern-1.5mu\mathsf{T}},$ (22)
where $\Lambda\in\mathbb{R}^{k\times k}$ is a diagonal matrix of real eigenvalues and $U\in\mathbb{R}^{k\times k}$ is the matrix of the corresponding orthogonal eigenvectors. Next, the numerical rank is determined based on the diagonal matrix $\Lambda$. A typical choice here is to truncate all eigenvalues $\lambda_{i}$ that are small in absolute value, e.g.,
$\lvert\lambda_{i}\rvert<k\cdot\max_{i=1,\dots,k}\lvert\lambda_{i}\rvert\cdot\epsilon,$
where $\epsilon$ is the machine precision. Note that the original matrix $D$ is assumed to be indefinite such that large positive as well as negative eigenvalues may occur, which need to be kept in the compression. For simplicity, assume that the eigenvalues in $\Lambda$ are ordered by absolute size such that we have
$\Lambda=\begin{bmatrix}\Lambda_{1}&0\\ 0&\Lambda_{2}\end{bmatrix}\quad\text{and}\quad U=\begin{bmatrix}U_{1}&U_{2}\end{bmatrix},$
where $\Lambda_{1},U_{1}$ contain the eigenvalues and eigenvectors that remain and $\Lambda_{2},U_{2}$ contain the eigenvalues and eigenvectors that are truncated. This results in the compression
$LDL^{\mkern-1.5mu\mathsf{T}}\approx QU_{1}\Lambda_{1}U_{1}^{\mkern-1.5mu\mathsf{T}}Q^{\mkern-1.5mu\mathsf{T}}=\widetilde{L}\widetilde{D}\widetilde{L}^{\mkern-1.5mu\mathsf{T}},$
where $\widetilde{L}=QU_{1}$ and $\widetilde{D}=\Lambda_{1}$, which is used to compute the solution to the matrix equations eqs. 15 and 16 instead of the original right-hand sides.